OpenStudy (anonymous):

check my work please? :) solve the system by substitution 2x - y + z = -4 z = 5 -2x + 3y - z = -10 I plugged it in and got (-8, 7, 5) is that correct?

OpenStudy (anonymous):

NEVER MIND I DID IT WRONG UGH I FORGOT THE NEGATIVES. one sec let me figure this out in the mean time please help!

OpenStudy (anonymous):

okay I changed it i do think it's (-8, -7, 5) because 2(-8) - (-7) + 5 = -4 -16 - (-7) + 5 = -4 -9 + 5 = -4 -4 = -4

OpenStudy (anonymous):

@Hero

hero (hero):

Just remember to test each equation. Just because it works for one doesn't mean it works for all.

OpenStudy (anonymous):

I did I plugged it into the other equation and got -10 = -10